[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Sep 24, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Sep 24 09:41:42 UTC 2017


HSU Shift Log for Sun Sep 24 2017 05:41:36 GMT-0400 (EDT)

Shift: A

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: None

AWIPS-2/NCP issues: None

Other Techincal issues: Website failed when special advisory was issued.  
Problem was related to using 0630 UTC as the issuing time. DAZ manually  
fixed problem on www01 server and restarted halo. Brief hiccup between WOC  
and Amazon Web Services caused a few storm table thumbnails and Maria  
archive to be missing/Forbidden on public site. Problem not observed on  
local server. Resolved itself after a few minutes.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170924/658d9e1c/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list